OpenStudy (anonymous):

Solve for x. 6e^2x = 7e^4x

OpenStudy (yoloshroom):

do you have multiple choice questions for this? and are we using E as a variable or as a number?

OpenStudy (luigi0210):

\(\Large 6e^{2x} = 7 e^{4x} \)?

OpenStudy (anonymous):

@noseboy908

OpenStudy (noseboy908):

Have you considered what @YoloShroom and @Luigi0210 have said?

OpenStudy (anonymous):

No multiple choice. e as in natural log. and waht @Luigi0210 is my problem.

OpenStudy (noseboy908):

Sorry, this is beyond me. Could you assist, @YoloShroom ?

OpenStudy (yoloshroom):

yeah let me take a look.

OpenStudy (luigi0210):

You're gonna need to take the logs of both sides, have you tried that?

OpenStudy (yoloshroom):

it could be x= 0, because the bases are the same, we can set the exponents to equal each other i believe. SO it may just be 2x=4x, then 2x=0, x=0.

OpenStudy (anonymous):

@Luigi0210 I haven't tried taking the logs. How would I do that? @YoloShroom How are the bases the same?

OpenStudy (yoloshroom):

you have the Natural log as both the bases, ^_^ May be wrong, just a suggestion.

OpenStudy (anonymous):

Oh, I was just confused cause I thought the 7 and the 6 were the bases. If the e's are the bases then they cancel each other out?

OpenStudy (luigi0210):

You can take the logs of both sides like this: \(\Large \color{red}{log}(6e^{2x} ) = \color{red}{log} (7e^{4x} ) \) Now you can apply log rules: \(\Large log(6)+log(e^{2x}) = log(7)+log(e^{4x} ) \) Make sense so far?

OpenStudy (yoloshroom):

OH GOD, i forgot the base rule.. my bad. Good call

OpenStudy (luigi0210):

Here's a hint: \(\large log(e^x) = x \)

OpenStudy (anonymous):

Okay. That makes sense.

OpenStudy (luigi0210):

You should end up with \(\Large log(6) +2x = log(7)+4x \) Get like terms on the same side: \(\Large -2x=log(7)-log(6)\) Now just divide both sides by -2 and you'll get your x \(\Large \ddot \smile \)

OpenStudy (anonymous):

I got x = -0.033

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

I think it's best to leave it as \(\Large x=-\frac{log(7)}{2}+\frac{log(6)}{2} \) But if you need a decimal then -0.0335 should work :)

OpenStudy (anonymous):

Okay. Thank you!

OpenStudy (luigi0210):

No problem :)